OpenStudy (anonymous):

What is the sum of an 8–term geometric series if the first term is 13 and the last term is –3,639,168?

OpenStudy (anonymous):

is the answer 4,366,999?

OpenStudy (anonymous):

if yes then you should rember that sum of the terms in GP is given by: \[S _{_{n}}=a(r ^{n}-1)/(r-1)\] if r>1 and \[S _{_{n}}=a(1-r ^{n})/(1-r)\] if r<1

OpenStudy (anonymous):

I got the answer...its this: –3,119,285...thanks for you help.